OpenStudy (anonymous):

It takes Sarah 5 days to paint a house. With the help of a friend they can paint the same house in 3 days. How fast can Sarah’s friend paint the house alone?

ganeshie8 (ganeshie8):

they both working together, take 3 days to complete the house => in 1 day, they complete 1/3rd of the house

ganeshie8 (ganeshie8):

lets say, her friend takes 'x' days to paint the house, ok

ganeshie8 (ganeshie8):

so, in one day her friend can complete 1/x of the house. and, in one day, Sarah can complete 1/5th of the house.

ganeshie8 (ganeshie8):

combining above 3 statements, 1/5 + 1/x = 1/3

ganeshie8 (ganeshie8):

you can solve for x now ?

OpenStudy (anonymous):

When I did it earlier I put everything into hours so that may be why I was confused.

ganeshie8 (ganeshie8):

ohh that may give us wierd numbers... hope u get the nub of solving it :)

ganeshie8 (ganeshie8):

now

OpenStudy (anonymous):

so when i do 1/5+1/x=1/3 would i multiply everything by 12 since its the LCD?

OpenStudy (anonymous):

by 15

OpenStudy (anonymous):

Im just confused as to what to do next.

ganeshie8 (ganeshie8):

thats right, we can multiply everything by LCD, but LCD here is not just 15, its 15x ! so multiply whole thing by 15x :)

ganeshie8 (ganeshie8):

15x(1/5 + 1/x) = 15x (1/3) see if u can simplify and solve x

OpenStudy (anonymous):

would if come out to be 3x+15=5x

ganeshie8 (ganeshie8):

Perfect !

ganeshie8 (ganeshie8):

solve it fully by subtracting 2x both sides,

OpenStudy (anonymous):

i got an answer of 7.5. would that be days or hours?

ganeshie8 (ganeshie8):

days. coz we worked it in days all the way till nw, good work :)

OpenStudy (anonymous):

ok. thanks alot!!!

ganeshie8 (ganeshie8):

np.. yw :)